# Modified scattering for the three dimensional Maxwell-Dirac system

### v4

@@ 1. Introduction and main results
The problem we will address in this work is the Cauchy problem for the Maxwell-Dirac system on the Minkowski space-time $\mathbb{R}^{1+3}$, which, expressed in the Lorenz gauge, has the form:
(1.1) $\left\{\begin{aligned} -i&\gamma^{\mu}\partial_{\mu}\psi+\psi=\gamma^{\mu}A_{\mu}\psi\\ &\Box A_{\mu}=-\overline{\psi}\gamma_{\mu}\psi\\ &\partial^{\mu}A_{\mu}=0.\end{aligned}\right$.
@@
  • We make no assumptions on the support of the initial data. Furthermore, we make very mild decay assumptions on the initial data at infinity. In particular, we use only three Lorentz vector fields in the analysis, which is close to optimal and significantly below anything that has been done before.
  • Rather than using arbitrarily high regularity, here we work with very limited regularity for the initial data, e.g. our three vector fields bound is simply in the energy space.
@@ 1.1. Previous work
Recent contributions to the massive (MKG) in $d=3$ include the work by Klainerman–Wang–Yang [32], establishing asymptotic properties for non-compactly supported data, and Fang–Wang–Yang [20], allowing for large Maxwell fields. Modified scattering behavior for the massive (MKG) in Lorenz gauge was established by Chen [11].
@@ 1.2. The Maxwell-Dirac system
with standard conventions for raising and lowering indices.
@@
$\mathscr{S}[A_{\mu},\psi]=\iint_{\mathbb{R}^{1+3}}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+i\langle\gamma^{\mu}\mathbf{D}_{\mu}\psi,\gamma^{0}\psi\rangle-\langle\psi,\psi\rangle\,dtdx.$
@@
(1.5) $\left\{\begin{aligned} &\partial^{\nu}F_{\mu\nu}=-\langle\psi,\gamma^{0}\gamma^{\mu}\psi\rangle\\ &i\gamma^{0}\gamma^{\mu}\mathbf{D}_{\mu}\psi=\gamma^{0}\psi.\end{aligned}\right$.
@@
A key feature of (1.5) is its invariance under gauge transformations meaning that given any solution $(A,\psi)$ of (1.5) and a real-valued function $\chi$, called gauge transformation, on $I\times\mathbb{R}^{3}$, the gauge transform $(\tilde{A},\tilde{\psi})=(A-\mbox{d}\chi,e^{i\chi}\psi)$ of $(A,\psi)$ is also a solution to (1.5). This in fact says that relative to this gauge transform we should think of a solution as being an equivalence class of functions that are solutions to our problem.
In order to address the well-posedness theory we need to remove the ambiguity arising from this invariance, for our system (1.5), and fix the gauge. Traditionally there are several gauges that have been used to address this issue. This includes for instance the Coulomb gauge $\partial_{j}A_{j}=0$, which leads to a mix of hyperbolic and elliptic equations, as well. Another possible gauge choice is the temporal gauge $A_{0}=0$, which retains causality but loses some ellipticity.
In our paper we impose the Lorenz gauge condition, which reads
@@
$\left\{\begin{aligned} -&i\gamma^{\mu}\partial_{\mu}\psi+\psi=\gamma^{\mu}A_{\mu}\psi\\ &\Box A_{\mu}=-\overline{\psi}\gamma_{\mu}\psi\\ &\partial^{\mu}A_{\mu}=0.\end{aligned}\right$.
@@
If one considers only the (self-contained) system formed by the first two equations in (1.1), then the initial data above can be chosen arbitrarily. However, if in addition one also adds the third equation, then the initial data is required to satisfy the following constraint equations
$\left\{\begin{aligned} &\dot{a}_{0}=\ \partial_{j}a_{j}\\ &\Delta a_{0}=\ \partial_{j}\dot{a}_{j}-|\psi_{0}|^{2},\end{aligned}\right$.
@@ 1.3. Functional spaces
    ${\bf E}:=\int\overline{\gamma^{j}D_{j}\psi}\psi+\bar{\psi}\psi+\frac{1}{2}|\nabla A|^{2}\,dx.\qquad$
In terms of terminology, our problem is called charge critical, and this is because the charge is measured in the critical space $L^{2}$. In $d=4$, the critical Sobolev space would change and the energy will be expressed in terms of these critical Sobolev spaces, leading to the terminology energy critical Maxwell-Dirac system.
@@ 1.4. Main results
  • we make the convention of using upper-case letters for multi-indices, e.g. $\partial^{I}_{x}=\partial_{x_{0}}^{i_{0}}\cdots\partial_{x_{d}}^{i_{d}}$ and $x^{I}={x_{0}}^{i_{0}}\cdots{x_{d}}^{i_{d}}$, where $I=(i_{0},\ldots,i_{d})$, and we write $I_{0}$ if $i_{0}=0$.
@@
(1.6) $\sum_{3|J_{0}|+|K_{0}|\leq 9}\|x^{J_{0}}\partial_{x}^{K_{0}}\psi_{0}\|_{L^{2}}+\|x^{J_{0}}\partial_{x}^{J_{0}+K_{0}}\psi_{0}\|_{L^{2}}+\|x^{J_{0}}\partial_{x}^{J_{0}+K_{0}}a\|_{\dot{H}^{\frac{1}{2}}}+\|x^{J_{0}}\partial_{x}^{J_{0}+K_{0}}\dot{a}\|_{\dot{H}^{-\frac{1}{2}}}\leq\varepsilon,$
@@
$\sum_{3|J|+|K|\leq 9}\|\Omega^{J}\partial^{K}\psi\|_{L^{2}}+\|\Omega^{J}\partial^{K}A\|_{\dot{H}^{\frac{1}{2}}}+\|\Omega^{J}\partial^{K}\partial_{t}A\|_{\dot{H}^{-\frac{1}{2}}}\leq\varepsilon t^{C\varepsilon},$
@@
$|\psi(t,x)|\lesssim\frac{\varepsilon}{{\langle t+|x|\rangle}^{\frac{3}{2}}\langle t\rangle^{\delta}}.$
@@
    Higher regularity: If the initial data for $(\psi,A)$ has additional regularity, then the hyperbolic space regularity of $(\rho^{\pm}_{\infty},a^{\pm}_{\infty})$ can be improved, as well as the decay rate for $\rho^{\pm}_{\infty}$ at the boundary of the unit ball. However, there is no improved decay rate for $a^{\pm}_{\infty}$; instead, $(1-v^{2})^{-\frac{1}{2}}a^{\pm}_{\infty}$ will always have a nondegenerate limit at the boundary.
@@
$|A(t,x)|\lesssim\ \frac{\varepsilon}{\langle t+|x|\rangle}\log\frac{2\langle t+r\rangle}{\langle t-r\rangle}.$
@@ Acknowledgments
The second author, M.I., was supported by the Sloan Foundation, by an NSF CAREER grant DMS-1845037, by a Simons Foundation via a Simons Fellowship, and by a Visiting Miller Professorship via the Miller Foundation. M.S. and S.H. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283. We would like to thank the Erwin Schrödinger International Institute in Vienna as well the University of California at Berkeley, where part of this paper has been written, for their hospitality.
@@ 3. Energy estimates for the linearized equation
(3.1) $\left\{\begin{aligned} -&i\gamma^{\mu}\partial_{\mu}\phi+\phi=\gamma^{\mu}A_{\mu}\phi+\gamma^{\mu}B_{\mu}\psi+F\\ &\Box B_{\mu}=-\overline{\phi}\gamma_{\mu}\psi-\overline{\psi}\gamma_{\mu}\phi+G_{\mu}\\ &\partial^{\mu}B_{\mu}=0.\end{aligned}\right$.
@@ 4.1. Energy estimates for vector fields
where $c\approx C_{0}$. Here we would like to use interpolation and Gronwall’s inequality as in the previous section. But since vector fields are time dependent, the interpolation should happen in a space-time setting. Because of this we will no longer be able to apply directly Gronwall’s inequality in time; instead it turns out that a dyadic time decomposition would address the issue; this is similar to the work of the second author in [25].
@@
    (4.7) $\|\nabla_{t,x}A\|_{L^{\infty}\dot{H}^{-\frac{1}{2}}(C_{T})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}L^{\frac{3}{2}}(C_{T})}.$
@@
(4.8) $\|A\|_{L^{\infty}L^{3}(C_{T})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}L^{\frac{3}{2}}(C_{T})}.$
@@
$\qquad+\int_{T}^{s}\|\partial_{t}\tilde{A}_{\lambda}(t)\|_{L^{2}}\|G_{\lambda}(t)\|_{L^{2}}\,dxdt,$
@@
$\|\tilde{A}(s)\|_{\dot{H}^{\frac{1}{2}}}^{2}+\|\partial_{t}\tilde{A}(s)\|_{\dot{H}^{-\frac{1}{2}}}^{2}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}^{2}+\|\partial_{t}\tilde{A}\|_{L^{\infty}\dot{H}^{-\frac{1}{2}}}\|\tilde{G}\|_{L^{1}\dot{H}^{-\frac{1}{2}}},$
where the space-time norms are taken over $T,s]\times{\mathbb{R}}^{3}$. Taking the supremum over $s\in[T,4T]$, an application of Young’s inequality and the embeddings $\dot{H}^{\frac{1}{2}}\hookrightarrow L^{3}$ and $L^{\frac{3}{2}}\hookrightarrow\dot{H}^{-\frac{1}{2}}$ yields
$\|\nabla_{t,x}A\|_{L^{\infty}\dot{H}^{-\frac{1}{2}}(C_{T})}$ $\lesssim\|\nabla_{t,x}\tilde{A}\|_{L^{\infty}\dot{H}^{-\frac{1}{2}}([T,4T]\times{\mathbb{R}}^{3})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|\tilde{G}\|_{L^{1}L^{\frac{3}{2}}([T,4T]\times{\mathbb{R}}^{3})}$
$\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}L^{\frac{3}{2}}(C_{T})}.$

### v5

@@ 1. Introduction and main results
The problem we will address in this work is the Cauchy problem for the Maxwell-Dirac system on the Minkowski space-time $\mathbb{R}^{1+3}$. In the Lorenz gauge
$\partial^{\mu}A_{\mu}=0,$
it has the form
(1.1) $\left\{\begin{aligned} -i&\gamma^{\mu}\partial_{\mu}\psi+\psi=\gamma^{\mu}A_{\mu}\psi,\\ &\Box A_{\mu}=-\overline{\psi}\gamma_{\mu}\psi.\end{aligned}\right$.
@@
  • We make no assumptions on the support of the initial data. Furthermore, we make very mild decay assumptions on the initial data at infinity. In particular, we use only three Lorentz vector fields in the analysis, which is close to optimal and significantly below all previous results on Maxwell-Dirac.
  • Rather than using arbitrarily high regularity, here we work with very limited regularity for the initial data. We only require energy-space bounds on three vector fields which are complemented with bounds on a maximum of nine derivatives without any further decay assumption.
  • To compare with the best results on the massive Maxwell-Klein-Gordon equation [32, 19], the global results require up to two derivatives with an additional decay of $r^{-\frac{5}{2}-}$ in the energy space, which is different from the one for Maxwell-Dirac.
@@
  • Similar to the hyperboloidal foliation method developed in [30, 48, 36, 37], we make essential use of the hyberbolic nature of the problem. In order to minimize the regularity and decay assumptions and to effectively employ the testing by wave packet method, we proceed differently in using space-time estimates in the $C_{TS}$ regions.
@@ 1.1. Previous work
Recent contributions to the massive (MKG) in $d=3$ include the work by Klainerman–Wang–Yang [32], establishing asymptotic properties for non-compactly supported data, and Fang–Wang–Yang [19], allowing for large Maxwell fields. A precise modified scattering result in the spirit of our Theorem 1.3 for the massive (MKG) in Lorenz gauge was established by Chen [10] and similarly for the wave-Klein-Gordon system by Chen–Lindblad in [11]. Both these systems and the Maxwell-Dirac system share the dispersive decay heuristics and the strong influence of the hyperbolic nature of the problem. We emphasize that in the Maxwell-Dirac case the Dirac component satisfies a first order equation, resulting in a different nonlinear coupling. Moreover, the intricate spinorial structure of the Maxwell-Dirac system requires new insights, as we will describe in further detail below Theorem 1.3.
@@ 1.2. The Maxwell-Dirac system
with standard conventions for raising and lowering indices, and $\Box=-\partial^{\nu}\partial_{\nu}$.
@@
$\mathscr{S}[A_{\mu},\psi]=\iint_{\mathbb{R}^{1+3}}\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+i\langle\gamma^{\mu}\mathbf{D}_{\mu}\psi,\gamma^{0}\psi\rangle-\langle\psi,\gamma^{0}\psi\rangle\,dtdx.$
@@
(1.5) $\left\{\begin{aligned} &\partial^{\nu}F_{\mu\nu}=\langle\psi,\gamma^{0}\gamma_{\mu}\psi\rangle\\ &i\gamma^{0}\gamma^{\mu}\mathbf{D}_{\mu}\psi=\gamma^{0}\psi.\end{aligned}\right$.
@@
A key feature of (1.5) is its invariance under gauge transformations meaning that given any solution $(A,\psi)$ of (1.5) and a real-valued function $\chi$, called gauge transformation, on $I\times\mathbb{R}^{3}$, the gauge transform $(\tilde{A},\tilde{\psi})=(A+\mbox{d}\chi,e^{i\chi}\psi)$ of $(A,\psi)$ is also a solution to (1.5). This in fact says that relative to this gauge transform we should think of a solution as being an equivalence class of functions that are solutions to our problem.
In order to address the well-posedness theory we need to remove the ambiguity arising from this invariance, for our system (1.5), and fix the gauge. Traditionally there are several gauges that have been used to address this issue. This includes for instance the Coulomb gauge $\partial_{j}A_{j}=0$, which leads to a mix of hyperbolic and elliptic equations. Another possible gauge choice is the temporal gauge $A_{0}=0$, which retains finite speed of propagation but loses some ellipticity. In our paper we impose the Lorenz gauge condition, which reads
@@
(1.6) $\left\{\begin{aligned} -i\gamma^{\mu}\partial_{\mu}\psi+\psi={}&\gamma^{\mu}A_{\mu}\psi\\ \Box A_{\mu}={}&-\overline{\psi}\gamma_{\mu}\psi\\ \partial^{\mu}A_{\mu}={}&0.\end{aligned}\right$.
@@
If one considers only the (self-contained) system formed by the two equations in (1.6), then the initial data above can be chosen arbitrarily. However, if in addition one also adds the third equation, then the initial data is required to satisfy the following constraint equations
(1.7) $\left\{\begin{aligned} \dot{a}_{0}={}&\ \partial_{j}a_{j}\\ \Delta a_{0}={}&\ \partial_{j}\dot{a}_{j}+|\psi_{0}|^{2},\end{aligned}\right$.
@@ 1.3. Functional spaces
    ${\bf E}:=\int i\overline{\gamma^{j}\mathbf{D}_{j}\psi}\psi+\overline{\psi}\psi+\frac{1}{2}(|E|^{2}+|B|^{2})\,dx,\qquad$
with the electric field $E=(F_{10},F_{20},F_{30})$ and the magnetic field $B=(F_{23},F_{31},F_{12})$. Note that $\bf E$} is not sign definite. In terms of terminology, our problem is called charge critical, and this is because the charge is measured in the critical space $L^{2}$. In $d=4$, the critical Sobolev space would change and the energy will be expressed in terms of these critical Sobolev spaces, leading to the terminology energy critical Maxwell-Dirac system.
@@ 1.4. Main results
  • We make the convention of using upper-case letters for multi-indices, e.g. $\partial^{I}=\partial_{x_{0}}^{i_{0}}\cdots\partial_{x_{d}}^{i_{d}}$ and $x^{I}={x_{0}}^{i_{0}}\cdots{x_{d}}^{i_{d}}$, where $I=(i_{0},\ldots,i_{d})$, and we write $I_{0}$ if $i_{0}=0$. We also use the shorthand notation $\partial^{k}=\{\partial^{I}\}_{|I|=k}$ and $\partial^{\leq k}=\{\partial^{I}\}_{|I|\leq k}$, for $k\in{\mathbb{N}}_{0}$. Similar as above, $k_{0}$ indicates that there are no $\partial_{t}(=\partial_{x_{0}})$ derivatives.
@@
(1.8) $\sum_{3j_{0}+k_{0}\leq 9}\||x|^{j_{0}}\partial^{\leq j_{0}+k_{0}}\psi_{0}\|_{L^{2}}+\||x|^{j_{0}}\partial^{j_{0}+k_{0}}a\|_{\dot{H}^{\frac{1}{2}}}+\||x|^{j_{0}}\partial^{j_{0}+k_{0}}\dot{a}\|_{\dot{H}^{-\frac{1}{2}}}\leq\varepsilon,$
@@
$\sum_{3|J|+|K|\leq 9}\|\Omega^{J}\partial^{K}\psi\|_{L^{2}}+\|\Omega^{J}\partial^{K}A\|_{\dot{H}^{\frac{1}{2}}}+\|\Omega^{J}\partial^{K}\partial_{t}A\|_{\dot{H}^{-\frac{1}{2}}}\lesssim\varepsilon\langle t\rangle^{C\varepsilon},$
@@
$|\psi(t,x)|\lesssim\frac{\varepsilon}{{\langle t+|x|\rangle}^{\frac{3}{2}}\langle|x|-t\rangle^{\delta}}.$
Concerning the Lorenz gauge condition, we point out that the compatibility conditions (1.7) between $a_{0}$ and $\psi_{0}$ give rise to the so called charge-problem, see Remark 1.5 below.
@@
    Higher regularity: If the initial data for $(\psi,A)$ has additional regularity, then the hyperbolic space regularity of $(\rho^{\pm}_{\infty},a^{\mu}_{\infty})$ can be improved, as well as the decay rate for $\rho^{\pm}_{\infty}$ at the boundary of the unit ball. However, there is no improved decay rate for $a^{\mu}_{\infty}$; instead, $(1-v^{2})^{-\frac{1}{2}}a^{\mu}_{\infty}$ will always have a limit at the boundary. This can be easily seen as a slight extension of the proof of Proposition 7.3, or it follows from the asymptotics of the fundamental solution for $(-1-\Delta_{H})$ in Remark 8.5. A similar result has been proven in the Maxwell-Klein-Gordon case in [10] and in the Wave-Klein-Gordon case in [11].
@@
$|A(t,x)|\lesssim\ \frac{\varepsilon}{\langle t+|x|\rangle}\log\frac{2\langle t+|x|\rangle}{\langle t-|x|\rangle}.$
@@
Remark 1.5.
The main results of this paper, Theorem 1.1 and Theorem 1.3, apply to the PDE system (1.1), with conditions on the full set of initial data $(a,\dot{a})$ and $\psi_{0}$. To obtain a Lorenz gauge solution to Maxwell-Dirac, the compatibility conditions (1.7) have to be satisfied. However, if $\psi_{0}\neq 0$, the equation $\Delta a_{0}=\partial_{j}\dot{a}_{j}+|\psi_{0}|^{2}$ shows that $a_{0}$ cannot decay faster than $1/|x|$, which is incompatible with the assumptions on $a_{0},\dot{a}_{0}$ in the above Theorems. This is known as the charge problem, see e.g. [9, Section 4] in the Maxwell-Klein-Gordon case. There is a standard solution to this, which we sketch here: With the conserved charge $\mathbf{q}=-\int|\psi_{0}|^{2}$ as above, define $Q(t,x)=\varrho(|x|-t)\frac{\mathbf{q}}{4\pi|x|}$ for a smooth function $\varrho$ with $\varrho(s)=1$ for $s\geq 1$ and $\varrho(s)=0$ for $s<1/2$. Then, the key property is that $\Box Q=0$ and this function is zero inside the light cone. The modified gauge field $\tilde{A}_{\mu}:=A_{\mu}-\delta_{0\mu}Q$ satisfies the same wave equations as $A_{\mu}$, but with modified initial data $(\tilde{a}_{\mu},\dot{\tilde{a}}_{\mu})$. Now, assume that $(a_{j},\dot{a}_{j},\psi_{0})$ satisfy the assumptions of Theorem 1.1 and $(a_{0},\dot{a}_{0})$ satisfy (1.7). Then,
$\tilde{a}_{0}=a_{0}-\varrho(|x|)\frac{\mathbf{q}}{4\pi|x|},\quad\dot{\tilde{a}}_{0}=a_{0}+\varrho^{\prime}(|x|)\frac{\mathbf{q}}{4\pi|x|}$
satisfies the assumptions. Note that $(\psi,\tilde{A}_{\mu})$ satisfies the modified system
(1.14) $-i\gamma^{\mu}\partial_{\mu}\psi$ $=\gamma^{\mu}\tilde{A}_{\mu}\psi+\gamma^{0}Q\psi,$
$\Box\tilde{A}_{\mu}$ $=-\overline{\psi}\gamma_{\mu}\psi.$
Then, the same conclusion of Theorem 1.1 holds true for $(\psi,\tilde{A}_{\mu})$. In particular, in view of the decay of $Q$, the same pointwise bounds hold for $(\psi,A_{\mu})$. The analysis remains unchanged, except for the estimates for $F$ in Section 4. To estimate the additional contribution in the Dirac equation, we use that $|\Omega^{J}\partial^{K}Q|\lesssim\varepsilon^{2}\langle t+|x|\rangle^{-1}$.
Similarly, the conclusion of Theorem 1.3 holds for the modified system. Here, the statement remains unchanged, as the system has not changed inside the light cone and we only use the bounds established in Theorem 1.1.
@@ Acknowledgments
The second author, M.I., was supported by the Sloan Foundation, by an NSF CAREER grant DMS-1845037, by a Simons Foundation via a Simons Fellowship, and by a Visiting Miller Professorship via the Miller Foundation. M.S. and S.H. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283. We would like to thank the Erwin Schrödinger International Institute in Vienna as well the University of California at Berkeley, where part of this paper has been written, for their hospitality. The authors would like to thank the anonymous referees for their careful remarks which helped a lot to improve the presentation.
@@ 3. Energy estimates for the linearized equation
(3.1) $\left\{\begin{aligned} -&i\gamma^{\mu}\partial_{\mu}\phi+\phi=\gamma^{\mu}A_{\mu}\phi+\gamma^{\mu}B_{\mu}\psi+F\\ &\Box B_{\mu}=-\overline{\phi}\gamma_{\mu}\psi-\overline{\psi}\gamma_{\mu}\phi+G_{\mu}.\end{aligned}\right$.
@@ 4.1. Energy estimates for vector fields
where $c\approx C_{0}$. Here we would like to use interpolation and Gronwall’s inequality as in the previous section. But since vector fields are time dependent, the interpolation should happen in a space-time setting. Because of this we will no longer be able to apply directly Gronwall’s inequality in time; instead it turns out that a dyadic time decomposition would address the issue; this is similar to the work of the second author in [25] and relies on ideas of Metcalfe-Tataru-Tohaneanu [41].
@@
    (4.7) $\|\nabla_{t,x}A\|_{L^{\infty}_{t}\dot{H}^{-\frac{1}{2}}_{x}(C_{T})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}(C_{T})}.$
@@
(4.8) $\|A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}(C_{T})}.$
@@
$\qquad+\int_{T}^{s}\|\partial_{t}\tilde{A}_{\lambda}(t)\|_{L^{2}}\|G_{\lambda}(t)\|_{L^{2}}\,dt,$
@@
$\|\tilde{A}(s)\|_{\dot{H}^{\frac{1}{2}}}^{2}+\|\partial_{t}\tilde{A}(s)\|_{\dot{H}^{-\frac{1}{2}}}^{2}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}^{2}+\|\partial_{t}\tilde{A}\|_{L^{\infty}_{t}\dot{H}^{-\frac{1}{2}}_{x}}\|\tilde{G}\|_{L^{1}_{t}\dot{H}^{-\frac{1}{2}}_{x}},$
where the space-time norms are taken over $T,s]\times{\mathbb{R}}^{3}$. Taking the supremum over $s\in[T,4T]$, an application of Young’s inequality and the embeddings $\dot{H}^{\frac{1}{2}}\hookrightarrow L^{3}$ and $L^{\frac{3}{2}}\hookrightarrow\dot{H}^{-\frac{1}{2}}$ yield
$\|\nabla_{t,x}A\|_{L^{\infty}_{t}\dot{H}^{-\frac{1}{2}}_{x}(C_{T})}$ $\lesssim\|\nabla_{t,x}\tilde{A}\|_{L^{\infty}_{t}\dot{H}^{-\frac{1}{2}}_{x}([T,4T]\times{\mathbb{R}}^{3})}\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|\tilde{G}\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}([T,4T]\times{\mathbb{R}}^{3})}$
$\lesssim\|\nabla_{t,x}A(T)\|_{\dot{H}^{-\frac{1}{2}}}+\|G\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}(C_{T})}.$